OpenStudy (anonymous):

Prove That Sin^0.5(x) * cos(x) - sin^(5/2)(x) cos(x) = cos^3(x)* squareroot(sinx) A better version in the post below

OpenStudy (anonymous):

\[\sin ^{.5} * \cos(x) - \sin ^{5/2}x * \cos(x) = \cos ^{3} * \sqrt{sinx}\]

myininaya (myininaya):

Did you try to factor sqrt(sin(x))*cos(x) out on the left hand side

myininaya (myininaya):

Do you know why I'm asking you to try that?

myininaya (myininaya):

That one side is written has a product and we need this side as a product. Factoring will do that. I notice they both have cos(x) and common and I know I need a sqrt(sin(x)) also because of the other side so I factored those items out. Let me know when you have tried this.

OpenStudy (anonymous):

How does one factor out sqrt(sinx) * cos(x) from sqrt(sin^5x) cosx....

myininaya (myininaya):

1/2+?=5/2

myininaya (myininaya):

\[u^\frac{1}{2} \cdot u^\frac{4}{2}=u^\frac{5}{2}\]

myininaya (myininaya):

Do you understand? I'm just using the law of exponents.

OpenStudy (anonymous):

Yes, it's confusing when one adds trig to it for some reason

myininaya (myininaya):

\[\sin^\frac{1}{2}(x)\cdot \sin^\frac{4}{2}(x)=\sin^\frac{5}{2}(x)\]

myininaya (myininaya):

4/2=2

OpenStudy (anonymous):

I understand that, can you just factor out the 2nd part? It becomes sqrt(sinx)*cosx (1-?...)

myininaya (myininaya):

so that still means you don't know that: \[\sin^\frac{1}{2}(x)\cdot \sin^\frac{4}{2}(x)=\sin^\frac{5}{2}(x)\] ?

myininaya (myininaya):

\[\sin^\frac{1}{2}(x)\cos(x)-\sin^{\frac{5}{2}}\cos(x)\] \[\sin^\frac{1}{2}(x)\cos(x)(1- ? )\] If you understand what I'm saying about 1/2+4/2=5/2 then you should be able to fill in that blank

OpenStudy (anonymous):

So it's \[(\sqrt{sinx} \times cosx) \times (1-\sin \frac{ 4 }{ 2 }) = \cos ^{3} \times \sqrt{sinx} ?\]

myininaya (myininaya):

well you are missing the x's in some parts and that 4/2 is an exponent

myininaya (myininaya):

and what is 1-sin^2(x)?

OpenStudy (anonymous):

cos^2 (x) is I remeber

myininaya (myininaya):

yep and cos(x)*cos^2(x) is the desired cos^3(x) that is on the other side of the equation

OpenStudy (anonymous):

so it's solved? :O

myininaya (myininaya):

no we proved both sides were the same we didn't solve

OpenStudy (anonymous):

well, that's what I meant, finding x values are a different story. (pi/4?) Thanks, I thought it would involving stating cosine in terms of sine, which would add more sqaure roots and throw me off further.